OpenStudy (ame2aw):

What volume of 6.00M HNO3 solution is needed to prepare 1500mL of 0.180M HNO3 solution?

OpenStudy (caozeyuan):

dilution formula is M1V1=M2V2

OpenStudy (caozeyuan):

M1=6,M2=.18,V2=1.5L

OpenStudy (rushwr):

@AME2AW do u still need help ?

OpenStudy (rushwr):

@caozeyuan Its actually C1* V1 = C2 * V2 C1 = initial concentration V1 = initial volume C2 = final concentration V2= final volume

OpenStudy (caozeyuan):

My Chem teacher use M as in molarity, because he says Concentration can be different units and only molarity is used in this equation

OpenStudy (ame2aw):

thank you for the responses, I now remember this section. and our teacher also used M but same formula.